1)     Given market demand Qd = 50 - P, and market supply P = Qs + 5

A) Find the market equilibrium price and quantity?

B) What would be the state of the market if market price was fixed at Birr 25 per unit?


1
Expert's answer
2022-01-16T13:31:27-0500

A) The market equilibrium price and quantity are:

Qd = Qs,

50 - P = P - 5,

2P = 55,

P = 27.5,

Q = 27.5 - 5 = 22.5 units.

B) If market price was fixed at Birr 25 per unit, then the shortage will occur as Qd > Qs.